OpenStudy (anonymous):

A company rents cars at $40 a day and 15 cents a mile. Its competitor's cars are $50 a day and 10 cents a mile. a. For each company, give a formula for the cost of renting a car for a day as a function of the distance traveled. b. On the same axes, graph both functions. c. Decide which company is cheaper. ***Not really sure how to start this off! @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

Start off by defining variables

jimthompson5910 (jim_thompson5910):

So how many variables will we need?

OpenStudy (anonymous):

umm 2? one for days and one for miles?

jimthompson5910 (jim_thompson5910):

close, but no

jimthompson5910 (jim_thompson5910):

one for miles, yes but you don't need one for days

jimthompson5910 (jim_thompson5910):

Let x = number of miles y = total cost (for the entire day)

OpenStudy (anonymous):

ohhh okay i see... so we need to come up with 2 equations right? one for company a and one for company b?

jimthompson5910 (jim_thompson5910):

For the first company, it says that it "rents cars at $40 a day and 15 cents a mile" So what is the equation?

OpenStudy (anonymous):

40x+15y ?

jimthompson5910 (jim_thompson5910):

what is y again?

OpenStudy (anonymous):

ohhh oops... so y=40+15x ?

jimthompson5910 (jim_thompson5910):

closer

jimthompson5910 (jim_thompson5910):

it's not $15 a mile though

OpenStudy (anonymous):

oh right! y=40+0.15x ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now find the other equation

OpenStudy (anonymous):

yay! okay would the other one be y=50+0.10x ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

awesome! :) so are we done with part a?

jimthompson5910 (jim_thompson5910):

yep, now onto part b

OpenStudy (anonymous):

okay, so that's graphing both of those equations?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

on the same axis

OpenStudy (anonymous):

okay, soooo for company one, it should hit the y axis at 40 right? but how do I graph the 0.15x part? :/

jimthompson5910 (jim_thompson5910):

0.15 is the slope 0.15 = 15/100 = 3/20

OpenStudy (anonymous):

ohh okay... so it'd be something like this?  did I graph that correctly?